OpenStudy (anonymous):

Write an equation of a line parallel to y = x -5 in slope-intercept form that passes through the point (-6, -1).

OpenStudy (jdoe0001):

what's the slope of -> y = x -5 anyway?

OpenStudy (anonymous):

y=mx+b= -1=1(-6)+b -1=-6+b 5=b y=x+5

OpenStudy (anonymous):

2/3